To determine freefall acceleration on a moon with no atmosphere, you drop your handkerchief off the roof of a baseball stadium there. The roof is 113 meters tall. The handkerchief reaches the ground in 18.2 seconds. What is freefall acceleration on this moon? (State the result as a positive quantity.)
1
Expert's answer
2010-11-16T07:13:14-0500
From the equations of the motion
h = gt2/2;
g = 2h/t2 = 2*113 / 18.22 = 0.68 m/s2